OpenStudy (anonymous):

explain the process of using graphing technology to solve a system of equations

OpenStudy (queelius):

The idea is this: we want a single point, (x, y), that satisfies both equations. Thus, we simply need to know which points (if any) are common to both functions, that is, where f(z) = g(z). A way to solve a system of equations inspired by this insight can be seen in the following: Given two equations, ax + by = c dx + ey = f Let's rewrite these equations as functions of x. ax + by = c => by = c - ax => y = f(x) = c/b - a/b*x dx + ey = f => ey = f - dx => y = g(x) = f/e - d/e*x We want the y's to be equal, thus we can solve for this: f(x) = g(x) => c/b - a/b*x = f/e - d/e*x. Solving for x, we get: d/e*x - a/b*x = f/e - c/b x(d/e - a/b) = f/e - c/b x = (f/e-c/b)/(d/e-a/b) Once we have x, we have y. (Note: y = c/b - a/b*x). For a linear system of equations, it is possible to get a single solution, no solution, or an infinite number of solutions. That happens if d/e-a/b = 0, or in other words, when they have the same slope. Visually, if you have the graph, you just need to see where the two lines intersect.